3/4x + 1 = x + 5/6 I know the answer is x=2/3 ut dont know how they got it.
Thanks

Answer by Alan3354(69443)

3/4x + 1 = x + 5/6
-----------------
Multiply thru by 24
18x + 24 = 24x + 20
Collect terms
4 = 6x
x = 2/3